OpenStudy (anonymous):

Find the vector v, and its magnitude is 5. v is orthogonal to x axis, and the dot product v.w = 6, and the vector w=i+2j.

OpenStudy (anonymous):

v.w=|v||w|cos theta

OpenStudy (amistre64):

a vetor that is perpendicular to the x axis with a magnitude of 5 is <0,5>

OpenStudy (amistre64):

is this vector in R^3?

myininaya (myininaya):

does v have ai+bj form or does v have ai+bj+ck form?

myininaya (myininaya):

2 dimensions or 3?

OpenStudy (amistre64):

if w=<1,2>....... v.w = (1.0)+(5.2) = 10...not 6

OpenStudy (amistre64):

the question makes no sense...

OpenStudy (amistre64):

(1.0)+(2.3) = 6 but then v =<0,3> with a magnitude of 3

OpenStudy (anonymous):

the answer is <0,3,4>

OpenStudy (amistre64):

the answer is...... clarify the question lol

myininaya (myininaya):

i got what amistre got

myininaya (myininaya):

a condradiction

OpenStudy (amistre64):

find the real roots of x^2 +9

OpenStudy (anonymous):

but i dont understand how he got the number 4 in the answer

OpenStudy (amistre64):

oh..by the way; when i say real rots i mean complex lol

myininaya (myininaya):

ok so we are looking at 3 dimensions not 2

OpenStudy (amistre64):

<1,2,0> <0,3,4> -------- 6...yes

OpenStudy (amistre64):

so....<0,3> is what it looks like when staring at it from the xy plane

OpenStudy (amistre64):

0^2 +3^2 +z^2 = |v|^2

OpenStudy (amistre64):

9+z^2 = 5^2 z^2 = 25 - 6 z = sqrt(16) = 4

OpenStudy (amistre64):

-6 means -9 :)

myininaya (myininaya):

thats what I did :)

OpenStudy (amistre64):

lol..its like haveing a parrot :)

OpenStudy (anonymous):

hey, may you type in the word and send me?

OpenStudy (amistre64):

just copy it off of here and put it in a text editor.. like SWord or notepad

OpenStudy (amistre64):

....my shift keys are on the fritz...

OpenStudy (anonymous):

Thanks a lot